if a variable accleration of -(1/2x^2) is working then at what time the particle will reach at x =0.25 m.if the velocity is 0 at x=1m at t=0.pls give me the analysed solution.

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Vijay,

Ans:- v dv/dx= -(1/2x^2)

integrating within the limits x=1 to x=x && v=0 to v=v we get,

v²=(1/x - 1)

or v=√(1/x - 1)

Again v= - dx/dt ( negatine sign comes, as with the increase of time x decreases) 

hence   - √x/√(1 - x) dx=dt

 Then integrating from x=1 to .25 and t=0 to t we get

t=∏/3 + √3/4 sec (ans)

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best vijay !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR